OpenStudy (anonymous):

Please help true or false ratios 7/9=56/72 10/13=140/180 16/9=192/108 10/1=1/10

OpenStudy (anonymous):

write each pair in lowest terms. 7/9 is already in lowest terms 56/72 has 8 as GCF - so divide 56 and 72 by GCF 56/72=7/9 the ratio 7/9=56/72 is correct

OpenStudy (anonymous):

10/13 is also already in lowest terms. 140/180 have 20 as GCF. so the 140/180=7/9 this ratio is false

OpenStudy (anonymous):

16/9 is already in lowest terms 192 and 108 have 4 as GCF. so 192/108=48/27. This ratio is false

OpenStudy (anonymous):

10/1 is equal to 10 1/10 is not equal to 10, so this ratio is also false

OpenStudy (anonymous):

only A is correct.

OpenStudy (anonymous):

Thank you